20 cm Х Frames: 43 Time: 0.1792 s Frame Rate: 240fps RESET LOCK 10 cm 20 cm 10 cm

Select one trial form video above ,make measurements that will allow you to predict the maximum height .Describe the process and make sure to include the measurements you made, your calculations,your prediction for how high the puck will go ,what assumptions you make.

please assist
Answer #1

given
distance travelled, d = 20 cm

= 0.20 m

time taken, t = 0.1792 s

so, speed of the puck v = 0.2/0.1792

= 1.116 m/s

let m is the mass of the puck.
let h is the maximum height reached by the puck.
Apply conservation of energy
final potential energy = initial kinetic energy

m*g*h = (1/2)*m*v^2

h = v^2/(2*g)

= 1.116^2/(2*9.8)

= 0.0635 m (or) 63.5 cm <<<<<<<<--------------Answer

I assumed the path is frictionless.